OpenStudy (anonymous):

Find the value of a1 for an infinite geometric series with s=9 and r=1/6

OpenStudy (jango_in_dtown):

S=a1/(1-r) so from the given values 9=a1/(1-1/6) implies a1=15/2

OpenStudy (jango_in_dtown):

S is the sum, r the common ratio and a1 the 1st term